# Regional Style and Color Transfer

### v2

@@ I Introduction
In the domain of computer vision, recent advancements in image segmentation techniques [1, 2, 3] have had a profound impact. These advancements facilitate the precise delineation of objects and regions within images. Traditional segmentation methods primarily relied on handcrafted features and heuristic algorithms. However, the emergence of deep learning-based approaches, particularly convolutional neural networks (CNNs) [4, 5, 6, 7, 8, 9], has revolutionized this field. CNN-based models like U-Net and Mask R-CNN [10, 11, 12, 13] have demonstrated exceptional performance in segmenting objects with intricate shapes and diverse appearances. This paves the way for more accurate and efficient image understanding tasks.
In parallel, The field of style transfer has concurrently witnessed significant progress in recent years, fueled by the growing interest in neural style transfer algorithms. These algorithms aim to apply the artistic style of a reference image to a target image while faithfully preserving content of the target image. Early methods employed optimization-based approaches. However, with the advent of deep learning, neural style transfer has evolved to leverage pre-trained CNNs like VGG and ResNet [14, 15, 16, 17], enabling the achievement of impressive results in real-time style transfer applications. These advancements have unveiled new avenues for creative expression and artistic rendering within the realm of digital media.
@@ II Methods
This section details the comprehensive approach adopted in this study to elucidate the influence of regionally applied stylistic and chromatic alterations. The proposed framework at fig 1 takes two input images: a content image and a style image. The model strategically partitions the style transfer process, applying the stylistic elements to the background region and the color transfer to the foreground region. To achieve this regional manipulation, semantic segmentation [19, 20] is leveraged to delineate the person foreground. Subsequently, the style-transferred background and color-transferred foreground are seamlessly integrated using a blending technique, effectively mitigating any potential artifacts at the boundary.

### v3

@@ I Introduction
In the domain of computer vision, recent advancements in image segmentation techniques [1, 2, 3, 4] have had a profound impact. These advancements facilitate the precise delineation of objects and regions within images. Traditional segmentation methods primarily relied on handcrafted features and heuristic algorithms. However, the emergence of deep learning-based approaches, particularly convolutional neural networks (CNNs) [5, 6, 7, 8, 9, 10, 11, 12], has revolutionized this field. CNN-based models like U-Net and Mask R-CNN [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23] have demonstrated exceptional performance in segmenting objects with intricate shapes and diverse appearances. This paves the way for more accurate and efficient image understanding tasks.
In parallel, The field of style transfer has concurrently witnessed significant progress in recent years, fueled by the growing interest in neural style transfer algorithms. These algorithms aim to apply the artistic style of a reference image to a target image while faithfully preserving content of the target image. Early methods employed optimization-based approaches. However, with the advent of deep learning, neural style transfer has evolved to leverage pre-trained CNNs like VGG and ResNet [24, 25, 26, 27, 28, 29, 30, 31, 32, 33], enabling the achievement of impressive results in real-time style transfer applications. These advancements have unveiled new avenues for creative expression and artistic rendering within the realm of digital media.
@@ II Methods
This section details the comprehensive approach adopted in this study to elucidate the influence of regionally applied stylistic and chromatic alterations. The proposed framework at fig 1 takes two input images: a content image and a style image. The model strategically partitions the style transfer process, applying the stylistic elements to the background region and the color transfer to the foreground region. To achieve this regional manipulation, semantic segmentation [35, 36, 37] is leveraged to delineate the person foreground. Subsequently, the style-transferred background and color-transferred foreground are seamlessly integrated using a blending technique, effectively mitigating any potential artifacts at the boundary.